OpenStudy (anonymous):

Physically what does it mean if the solution of quadratic equation are complex numbers?

OpenStudy (unklerhaukus):

so the roots are not real, this means that the curve does not cut the x axis

OpenStudy (anonymous):

is it it have no solution

OpenStudy (unklerhaukus):

complex solutions are solutions

OpenStudy (unklerhaukus):

for example \[x^2+1=0\] \[x=\frac{-0\pm\sqrt{0^2-4(1)(1)}}{2}\]\[x=\frac{\pm2i}2=\pm i\] if we plot \[y(x)=x^2+1=0\]